Consider a 30-year mortgage at an interest rate of 9% compounded monthly with a $1300 monthly payment. What is the total amount paid in interest?
Answer #1

The present value of the mortgage is first computed here as:

1300 1300

This is a sum of 360 terms with common ratio as 1/(1 + 0.09/12) = 1/1.0075

Therefore the sum of geometric progression here is computed as:

= 1300 + I-10-60) *(1- · = 162778.1736 )

Total amount paid is computed here as: = 1300*360 = 468000

Therefore the total amount paid in interest is computed here as:

= Total amount paid - Total initial Principal payment

= 468,000 - 162,778.1736 = 305221.8264

Therefore 305221.8264 is the required amount here.